# Fast and Accurate Decoder for the XZZX code Using Simulated Annealing

### v1

@@ V.1 Simulation setting
In the Monte Carlo simulation, we generate samples of actual error configurations $C$ by means of pseudo-random numbers, drawn from the physical error rate $p$ and the error bias $p_{x}:p_{y}:p_{z}$ we specify. Performing the stabilizer measurement on the actual error configurations $C$ yields the syndrome $S$. Then, we execute each decoder to obtain the recovery configuration $\bar{C}$. If the product $C\,\bar{C}$ belongs to the (non-trivial) logical operator class $L_{X}],[L_{Y}]$ or $L_{Z}]$, then the logical error occurs. By repeating the aforementioned procedure $N_{\mathrm{sample}}$ times, we estimate the logical error rate $P_{\mathrm{L}}$. Except in Sec. V.4, we employ $N_{\mathrm{sample}}=10^{6}$ to demand sufficient statistical precision. Except in Sec. V.3, we refer to Algorithm 7 as ‘our SA decoder’ and employ the number of inverse temperatures $N_{\beta}=100$.
Although our SA decoder can be applied to any type of biased noise, we focus on $Y$-biased noise such that $X$ and $Z$ errors occur with the same probabilities $p_{x}=p_{z}$, which is the simplest case where the MWPM decoder cannot produce a correct recovery configuration. By the condition $p_{x}=p_{z}$, we can express as $p_{x}:p_{y}:p_{z}=1:\eta:1$ where $\eta$ indicates the $Y$-biased factor. Since the depolarizing noise is the case where $p_{x}=p_{y}=p_{z}=1/3$, it corresponds to $p_{x}:p_{y}:p_{z}=1:1:1$, i.e., $\eta=1$. The more $Y$-biased, the less accurate the MWPM decoder.
@@ V.2 Logical error rates
Let us compare the decoding accuracy of our SA decoder described in Algorithm 7 and the CPLEX decoder described in Sec. A. First, we examine how the logical error rate varies with the code distance $d$.
Fig. 3 depicts the logical error rates $P_{\mathrm{L}}$ under the depolarizing noise and for the code distances $d=5,7,9$. In Fig. 3, for every $d$, our SA decoder achieves CPLEX’s rates, which is optimal among MLE-type decoders. In lower physical error rates $p$, our SA decoder outperforms the CPLEX decoder. This is because, unlike the CPLEX decoder, our SA decoder accounts for the degeneracy of error configurations by setting the final temperature to the Nishimori temperature $\beta_{\mathrm{nishimori}}$ in Eq. (12).
@@
For both the SA and CPLEX decoders, from the intersection point of these curves, we determine the threshold $p_{\mathrm{th}}\simeq 0.18$, which approximately equal to $p_{\mathrm{th}}=0.187$ obtained by the MPS decoder [9]. Below the threshold $p_{\mathrm{th}}$, as $d$ increases, the logical error rate decreases, as expected.
@@
Figure 3: Logical error rates under the depolarizing noise and the code distance $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder for all $d$. We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy. The error bars indicate one standard deviation.
@@
Fig. 4 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. In Fig. 4, for every $d$, our SA decoder achieves CPLEX’s rates, which are optimal. For both SA and CPLEX decoders, the intersection point of these curves lies around $p=0.2$, which we consider to be the threshold. Below this threshold, the logical error rate decreases as $d$ increases.
@@
Figure 4: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and for the code distance $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder. We employ $N_{\mathrm{SA}}=100$, which suffices to attain CPLEX’s accuracy. The error bars indicate one standard deviation.
@@
Fig. 5 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. In Fig. 5, no intersection point is observed for $d=5,7,9$; hence, we infer that the threshold lies beyond the plotted range, i.e., $p_{\mathrm{th}}>0.2$. On the other hand, for $d=9$ in Fig. 5, our SA decoder does not achieve CPLEX’s rate. The curves for $d=5,7,9$ do not intersect, due to the critical slowing down of the MCMC method. For our SA decoder, the curves for $d=7,9$ intersect around $p=0.11$, which we regarded as the threshold. When $p<0.11$, the logical error rate decreases as $d$ increases.
@@
Figure 5: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and for the code distance $d=5,7,9$. Our SA decoder cannot achieve the logical error rate of the CPLEX decoder for $d=9$, even by $N_{\mathrm{SA}}=100$. The error bars indicate one standard deviation.
@@
Next, we investigate the dependence of the logical error rates $P_{\mathrm{L}}$ on the number of SA, $N_{\mathrm{SA}}$ for fixed code distance $d=9$. For comparison, we also calculate the rates by our greedy matching decoder, i.e., the GREEDY_MATCHING function described in Algorithm 5 and the MWPM decoder i.e., the MWPM function described in Algorithm 4.
@@
Figure 6: Logical error rates under the depolarizing noise and $N_{\mathrm{SA}}=1,\dots,10$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
Figure 7: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and $N_{\mathrm{SA}}=1,\dots,100$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
Figure 8: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and $N_{\mathrm{SA}}=1,\dots,100$. We employ the code distance $d=9$. The error bars indicate one standard deviation.
In Figs. 68, we observe that the logical error rate $P_{\mathrm{L}}$ of our SA decoder monotonically decreases as $N_{\mathrm{SA}}$ increases. For $N_{\mathrm{SA}}\geq 2$, our SA decoder achieves logical error rates lower than our greedy matching decoder and is therefore an improvement over our greedy matching. For $N_{\mathrm{SA}}\geq 3$, our SA decoder is more accurate than the MWPM decoder.
To achieve the accuracy by the CPLEX decoder, under the depolarizing noise in Fig. 6, $N_{\mathrm{SA}}=10$ is sufficient. Although under the noise $p_{x}:p_{y}:p_{z}=1:5:1$ in Fig. 7, $N_{\mathrm{SA}}=100$ is almost sufficient. Under the noise $p_{x}:p_{y}:p_{z}=1:10:1$ in Fig. 8, $N_{\mathrm{SA}}=100$ is not sufficient. Hence, to attain the optimal logical error rate, we observe that the more strongly $Y$-biased noise, the more $N_{\mathrm{SA}}$ our SA decoders need.
V.3 Convergence of the SA decoders
To construct a pure error, there are the following three methods:
@@ V.3 Convergence of the SA decoders
Every incorrect syndrome is directly connected with closer one in the top and bottom boundaries for white faces (the left and right boundaries for gray faces). We employ the SA decoder described in Algorithm 1.
@@
It is Algorithm 5 invoked in the SA decoder of Algorithm 1, described in Sec. IV.3. For every invocation, it produces the same pure error.
@@
It is Algorithm 6 invoked in the SA decoder of Algorithm 7, described in Sec. IV.4. For every invocation, it may produce a different pure error. We use this method in the present paper, unless otherwise stated.
@@
where $P^{\mathrm{(SA)}}_{\mathrm{L}}$ and $P^{\mathrm{(CPLEX)}}_{\mathrm{L}}$ are the logical error rates obtained by the SA and CPLEX decoders, respectively. Here, the power $2/(d+1)$ is to cancel out $d$-dependence. As for the ratio $R_{\mathrm{L}}$, the lower value implies that our SA decoder is more accurate than the CPLEX decoder, and the value $1$ indicates that our SA decoder attains CPLEX decoder’s accuracy.
Fig. 9 shows the ratio $R_{\mathrm{L}}$ for varying the number of inverse temperatures $N_{\beta}$ under the depolarizing noise. Similarly, Fig. 10 and Fig. 11 show the ratio $R_{\mathrm{L}}$ under the $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. In Figs. 911, the ratio values $R_{\mathrm{L}}$ obtained by the greedy matching and MWPM decoders are drawn by dotted gray and dashed-dotted brown horizontal lines, respectively.
@@
Figure 9: Convergence of our SA decoder for different construction methods of pure errors under the depolarizing noise. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
@@
Figure 10: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:5:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
@@
Figure 11: Convergence of our SA decoder for different construction methods of pure errors under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:10:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$.
In Figs. 911, under all the noises and for all the pure error construction methods, we can see that the ratio $R_{\mathrm{L}}$ broadly decreases to the value $1$ as $N_{\beta}$ increases. For $N_{\beta}\geq 20$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods outperform the MWPM decoder.
Setting $N_{\beta}=0$ disables SA. Hence, $N_{\mathrm{SA}}$ indicates the number of not SAs but pure errors. In this setting, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ decoders become their MLD variants. In all the cases (a), (b), (c) of Figs. 911, the ‘Greedy matching (different)’ method with $N_{\beta}=0$ achieves higher accuracy than the MWPM decoder.
For all values of $N_{\beta}$ in Figs. 911, ‘Greedy matching (different)’ is the highest accurate, followed by ‘Greedy matching (same)’, while the ‘Boundary’ method is the least accurate. The causes of these accuracy differences are considered as follows:
  • Whereas the ‘Boundary’ method ignores the actual error probabilities $p_{x}$, $p_{y}$, $p_{z}$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods take $p_{x}$, $p_{z}$ (but not $p_{y}$) into account. Consequently, the initial error configuration produced by ‘Greedy matching (same)’ is typically more similar to the actual error configuration than that produced by the ‘Boundary’ method. Thus, the ‘Greedy matching (same)’ method is more accurate than the ‘Boundary’ method.
  • The ‘Greedy matching (different)’ method may produce a different initial error configuration for each SA. It leads to numerous variations of initial configurations, increasing the chance of reaching lower-energy states and thereby reducing the logical error rate.
@@ V.4 Comparison of decoding time with other decoding algorithms
First, Fig. 12 shows the elapsed times for the SA and CPLEX decoders for various code distances $d$ under the depolarizing noise. Similarly, Figs. 13 and 14 show the elapsed times under $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. For measuring the elapsed times, we execute the SA and CPLEX decoders without any parallelization. From Figs. 1214, we can see that the larger $d$, the longer time both the SA and CPLEX decoders take.
@@
Figure 12: Elapsed times for the SA and CPLEX decoders under the depolarizing noise. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
@@
Figure 13: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
@@
Figure 14: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. The vertical axis is logarithmic scale. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (a) and (b), the dashed orange line implies elapsed time divided by $4N_{\mathrm{SA}}$, which is anticipated time assuming parallel execution of every SA. In (c), the CPLEX decoder indicates exponential time for the code distance $d$.
For each decoder, we examine the dependence on $d$ and $p$. For all the physical error rates $p$, our SA decoder requires the same time, because the computational cost is determined solely by $N_{\beta}$ and $d$. Recalling that for each inverse temperature, we perform the Metropolis step $\mathcal{O}(N_{\mathrm{stabilizer}})=\mathcal{O}(d^{2})$ times, the total elapsed time scales accordingly. On the other hand, the CPLEX decoder exhibits exponential scaling in computation time with respect to the code distance $d$. This exponentiality is clearly visible for the case (c) $p=0.15$ in Figs. 1214, where it is prohibitive to calculate for $d>11$.
Let us compare the elapsed times of our SA decoder and the CPLEX decoder. For $p=0.15$, our SA decoder is clearly faster than the CPLEX decoder, whereas for $p=0.02,0.1$, it requires more time than the CPLEX decoder.
In each of Figs. 1214, ‘SA (divided by $4N_{\mathrm{SA}}$)’ drawn with the dashed orange line is the elapsed time divided by $4N_{\mathrm{SA}}=40$, indicating anticipated time in parallelized execution. By comparing the dashed orange line and the blue line, we can see that our SA decoder is faster than the CPLEX decoder. Consequently, if our SA decoder is parallelized, it outperforms the CPLEX decoder for all the values $d$ and $p$ we employ.

### v2

@@ V.1 Simulation setting
In Monte Carlo simulations, we generate samples of true error configurations $\bar{C}$ by means of pseudo-random numbers, drawn from the physical error rate $p$ and the error bias $p_{x}:p_{y}:p_{z}$ we specify. Performing the stabilizer measurement on the true error configurations $\bar{C}$ yields the syndrome $S$. Then, we execute each decoder to obtain the recovery configuration $C$. If the product $\bar{C}\,C$ belongs to a nontrivial logical operator class, i.e., if $\bar{C}\,C]\in\{[L_{X}],[L_{Y}],[L_{Z}]\}$, then the logical error is said to occur. By repeating the aforementioned procedure $N_{\mathrm{sample}}$ times, we estimate the logical error rate $P_{\mathrm{L}}$. Except in Sec. V.4, we employ $N_{\mathrm{sample}}=10^{6}$ to ensure sufficient statistical precision. Except in Sec. V.3, we refer to Algorithm 7 as our SA decoder and employ the number of inverse temperatures $N_{\beta}=100$.
Our SA decoder is applicable to arbitrary biased Pauli noise provided that the decoder is supplied with the error probabilities $p_{x},p_{y},p_{z}$ used to define $\mathcal{H}(C)$. We focus on $Y$-biased noise as a stress test for the MWPM decoder, which neglects $Y$-induced correlations. For simplicity, we consider the symmetric case $p_{x}=p_{z}$. Under the condition $p_{x}=p_{z}$, we can write the bias as $p_{x}:p_{y}:p_{z}=1:\eta:1$ where $\eta$ denotes the $Y$-biased factor. Since the depolarizing noise is the case where $p_{x}=p_{y}=p_{z}=p/3$, it corresponds to $p_{x}:p_{y}:p_{z}=1:1:1$, i.e., $\eta=1$.
@@ V.2 Logical error rates
Let us compare the decoding accuracy of our SA decoder described in Algorithm 7 with the CPLEX decoder described in Appendix A. The CPLEX decoder serves as an optimal minimum-energy (MAP-configuration) decoder for the objective $\mathcal{H}(C)$ under the constraint $\partial C=S$. First, we examine how the logical error rate varies with the code distance $d$.
Fig. 3 depicts the logical error rates $P_{\mathrm{L}}$ under the depolarizing noise for the code distances $d=5,7,9$. In Fig. 3, our SA decoder achieves the logical error rates of the CPLEX decoder within statistical uncertainty for all $d$ shown.
@@
For both the SA and CPLEX decoders, from the crossing of the finite-distance curves, we estimate a finite-size threshold $p_{\mathrm{th}}\simeq 0.18$. This estimate is close to the previously reported value $p_{\mathrm{th}}=0.187$ obtained using an MPS decoder [xzzx_nature]. Below the threshold $p_{\mathrm{th}}$, the logical error rate decreases with increasing $d$, as expected.
@@
Figure 3: Logical error rates under the depolarizing noise for the code distances $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder for all $d$. We employ $N_{\mathrm{SA}}=100$, which suffices to attain the logical error rates of the CPLEX decoder. The error bars indicate one standard deviation.
@@
Fig. 4 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. In Fig. 4, for every $d$, our SA decoder achieves the logical error rates of the CPLEX decoder, which are optimal. For both SA and CPLEX decoders, the intersection point of these curves lies around $p\simeq 0.2$, which we consider to be a finite-size threshold estimate. Below this threshold, the logical error rate decreases as $d$ increases.
@@
Figure 4: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ for the code distances $d=5,7,9$. Our SA decoder achieves the logical error rates of the CPLEX decoder. We employ $N_{\mathrm{SA}}=100$, which suffices to attain the logical error rates of the CPLEX decoder. The error bars indicate one standard deviation.
@@
Fig. 5 shows the logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. In Fig. 5, no intersection point for the CPLEX decoder is observed for $d=5,7,9$; hence, we expect that the threshold for the CPLEX decoder lies beyond the plotted range, i.e., $p_{\mathrm{th}}>0.2$. On the other hand, for $d=9$ in Fig. 5, our SA decoder does not achieve the logical error rates of the CPLEX decoder. The lack of a clear crossing for $d=5,7,9$ may be related to slow mixing of the MCMC method. For our SA decoder, the curves for $d=7,9$ intersect around $p=0.11$, which we take as an estimated finite-size threshold. When $p<0.11$, the logical error rate decreases as $d$ increases.
@@
Figure 5: Logical error rates under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ for the code distances $d=5,7,9$. Our SA decoder cannot achieve the logical error rate of the CPLEX decoder for $d=9$, even by $N_{\mathrm{SA}}=100$. The error bars indicate one standard deviation.
@@
Next, we investigate the dependence of the logical error rates $P_{\mathrm{L}}$ on the number of SA runs, $N_{\mathrm{SA}}$ for fixed code distance $d=9$. For comparison, we also calculate the rates by our greedy matching decoder, i.e., the greedy_matching function described in Algorithm 5 and the MWPM decoder i.e., the MWPM function described in Algorithm 4.
@@
Figure 6: Logical error rates under the depolarizing noise and $N_{\mathrm{SA}}=1,\dots,10$ for the code distance $d=9$. The error bars indicate one standard deviation.
Figure 7: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$ and $N_{\mathrm{SA}}=1,\dots,100$ for the code distance $d=9$. The error bars indicate one standard deviation.
Figure 8: Logical error rates under $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$ and $N_{\mathrm{SA}}=1,\dots,100$ for the code distance $d=9$. The error bars indicate one standard deviation.
In Figs. 68, we observe that the logical error rate $P_{\mathrm{L}}$ of our SA decoder decreases monotonically as $N_{\mathrm{SA}}$ increases. For $N_{\mathrm{SA}}\geq 2$, our SA decoder achieves logical error rates lower than our greedy matching decoder and is therefore an improvement over our greedy matching. For $N_{\mathrm{SA}}\geq 3$, our SA decoder achieves a lower logical error rate than the MWPM decoder.
To achieve the accuracy by the CPLEX decoder, under the depolarizing noise in Fig. 6, $N_{\mathrm{SA}}=10$ is sufficient. Although under the noise $p_{x}:p_{y}:p_{z}=1:5:1$ in Fig. 7, $N_{\mathrm{SA}}=100$ is almost sufficient. Under the noise $p_{x}:p_{y}:p_{z}=1:10:1$ in Fig. 8, $N_{\mathrm{SA}}=100$ is not sufficient. Hence, to attain the optimal logical error rate, we observe that the more strongly $Y$-biased the noise is, the larger $N_{\mathrm{SA}}$ our SA decoder needs.
V.3 Effect of initial error configuration on SA performance
To construct a reference configuration $R(S)$, we consider the following three methods:
@@ V.3 Effect of initial error configuration on SA performance
Each syndrome defect is connected to the nearer of the two boundaries (top or bottom for white faces; left or right for gray faces). We call this reference configuration construction routine, instead of the greedy_matching function in Algorithm 1.
@@
We use Algorithm 5 as the initializer in Algorithm 1; it returns the same reference configuration on every call.
@@
It is Algorithm 6 invoked in the SA decoder of Algorithm 7, described in Sec. IV.4. For every invocation, it may produce a different reference configuration. We use this method in this paper, unless otherwise stated.
@@
where $P^{\mathrm{(SA)}}_{\mathrm{L}}$ and $P^{\mathrm{(CPLEX)}}_{\mathrm{L}}$ are the logical error rates obtained by the SA and CPLEX decoders, respectively. Here, the exponent $2/(d+1)$ is introduced to cancel out $d$-dependence. A value $R_{\mathrm{L}}=1$ indicates that the SA decoder matches the CPLEX benchmark. Values above $1$ indicate worse performance than CPLEX.
Fig. 9 shows the ratio $R_{\mathrm{L}}$ for varying the number of inverse temperatures $N_{\beta}$ under the depolarizing noise. Similarly, Fig. 10 and Fig. 11 show the ratio $R_{\mathrm{L}}$ under the $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. In Figs. 911, the ratio values $R_{\mathrm{L}}$ obtained by the greedy matching and MWPM decoders are drawn by dotted gray and dash-dotted brown horizontal lines, respectively.
@@
Figure 9: Convergence of our SA decoder for different construction methods of reference configurations under the depolarizing noise. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$. Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
@@
Figure 10: Convergence of our SA decoder for different construction methods of reference configurations under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:5:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$. Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
@@
Figure 11: Convergence of our SA decoder for different construction methods of reference configurations under the $Y$-biased noise such that $p_{x}:p_{y}:p_{z}=1:10:1$. The horizontal axis is the number of inverse temperatures $N_{\beta}$, and the vertical axis is the ratio $R_{\mathrm{L}}$. We employ $d=9$ and $N_{\mathrm{SA}}=10$. (a) $p=0.04$ (b) $p=0.1$ (c) $p=0.15$. Dotted gray and dash-dotted brown horizontal lines show the ratios obtained by the greedy matching and MWPM decoders, respectively.
In Figs. 911, under all the noises and for all the reference configuration construction methods, we see that the ratio $R_{\mathrm{L}}$ broadly decreases to the value $1$ as $N_{\beta}$ increases. For $N_{\beta}\geq 20$, the ‘Greedy matching (same)’ and ‘Greedy matching (different)’ methods outperform the MWPM decoder.
For the ‘Greedy matching (same)’ and ‘Greedy matching (different)’, when $N_{\beta}=0$, the simulated_annealing function performs no Metropolis updates, so the decoder compares only the energies of the initial configurations. We refer to these variants as “initial-configuration-only” variants. In all the cases (a), (b), (c) of Figs. 911, the initial-configuration-only variant of the ‘Greedy matching (different)’ method achieves higher accuracy than the MWPM decoder.
In Figs. 911, for all plotted values of $N_{\beta}$ and for all noise biases, ‘Greedy matching (different)’ consistently yields the lowest $R_{\mathrm{L}}$ among the three methods, followed by ‘Greedy matching (same)’, while the ‘Boundary’ method performs worst. We attribute these accuracy differences to the following factors:
  • The ‘Boundary’ method ignores the error probabilities $p_{x}$, $p_{y}$, and $p_{z}$, whereas ‘Greedy matching (same)’ incorporates $p_{x}$ and $p_{z}$ through the weights $\tilde{\alpha}_{x}$ and $\tilde{\alpha}_{z}$. As a result, ‘Greedy matching (same)’ yields a bias-informed starting configuration.
  • The ‘Greedy matching (different)’ method randomizes the tie-breaking among equal-distance pairs, generating diverse initial configurations across runs. This diversity increases the chance of reaching lower-energy configurations and thereby reduces the logical error rate.
@@ V.4 Comparison of decoding time with other decoding algorithms
First, Fig. 12 shows the elapsed times for the SA and CPLEX decoders for various code distances $d$ under the depolarizing noise. Similarly, Figs. 13 and 14 show the elapsed times under $Y$-biased noises where $p_{x}:p_{y}:p_{z}=1:5:1$ and $1:10:1$, respectively. For measuring the elapsed times, we execute the SA and CPLEX decoders without any parallelization. From Figs. 1214, we see that the larger $d$ is, the longer both the SA and CPLEX decoders take.
@@
Figure 12: Elapsed times for the SA and CPLEX decoders under the depolarizing noise. The vertical axis denotes elapsed time with logarithmic scale. The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
@@
Figure 13: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:5:1$. The vertical axis is on a logarithmic scale. The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
@@
Figure 14: Elapsed times for the SA and CPLEX decoders under the $Y$-biased noise where $p_{x}:p_{y}:p_{z}=1:10:1$. The vertical axis is on a logarithmic scale. The orange dashed line shows the elapsed time of SA divided by $4N_{\mathrm{SA}}$, an idealized runtime estimate assuming perfect parallel efficiency across $4N_{\mathrm{SA}}$ independent SA instances and negligible overhead. (a) $p=0.02$ (b) $p=0.1$ (c) $p=0.15$. In (c), the CPLEX decoder exhibits exponential scaling for the code distance $d$.
For each decoder, we examine the dependence on $d$ and $p$. For all the physical error rates $p$, our SA decoder takes approximately the same time, because the iteration counts are determined solely by $N_{\beta}$ and $d$ and does not depend on the syndrome weight $|S|$. Recalling that for each inverse temperature, we perform the Metropolis step $\mathcal{O}(N_{\mathrm{stabilizer}})=\mathcal{O}(d^{2})$ times, the total elapsed time scales accordingly. On the other hand, the CPLEX decoder exhibits exponential scaling in computation time with respect to the code distance $d$. This exponential behavior is clearly visible for the case (c) $p=0.15$ in Figs. 1214, where it is prohibitive to calculate for $d>11$.
Let us now compare the elapsed times of our SA decoder and the CPLEX decoder. For $p=0.15$, our SA decoder is clearly faster than the CPLEX decoder, whereas for $p=0.02,0.1$, it requires more time than the CPLEX decoder.
In each of Figs. 1214, ‘SA (divided by $4N_{\mathrm{SA}}$)’ drawn with the orange dashed line shows the elapsed time divided by $4N_{\mathrm{SA}}=40$, which corresponds to an idealized runtime assuming perfect parallel efficiency (i.e., negligible overhead and ideal strong scaling). Comparing the idealized parallel estimate (orange dashed) with the CPLEX time (blue) suggests that, under near-perfect scaling across the $4N_{\mathrm{SA}}$ independent runs, SA could be faster than CPLEX over the tested $(d,p)$ range. Quantifying the achievable parallel efficiency and overhead is left for future work.